ganeshie8 (ganeshie8):

Find one prime divisor of \(2^{5903}-1\)

ganeshie8 (ganeshie8):

if it helps use wolfram/programming/anything

OpenStudy (fibonaccichick666):

ah, nope

OpenStudy (fibonaccichick666):

I'm gonna go with 5903 itself. Since 5903 is prime and that makes thins one of thos eqs from number theory

ganeshie8 (ganeshie8):

dividing by 5903 leaves a remainder 1

OpenStudy (fibonaccichick666):

I think it goes like \[2^p-1\equiv 0modp\]

OpenStudy (fibonaccichick666):

hmmm

ganeshie8 (ganeshie8):

\[2^{p-1}-1\equiv 0 \pmod p\] is true by fermat

OpenStudy (fibonaccichick666):

p-1... darn

ganeshie8 (ganeshie8):

but i think you're on right track

ganeshie8 (ganeshie8):

idk maybe not.. my method is bit different...

OpenStudy (fibonaccichick666):

idk, I was guessing cause it reminded me of that haha

OpenStudy (fibonaccichick666):

hmm well, i'm gonna say that it is whatever \[2^{5903}-1\] is since 5903 is prime

OpenStudy (fibonaccichick666):

so that number must be prime too

ganeshie8 (ganeshie8):

it is composite

ganeshie8 (ganeshie8):

i had to find a divisor in proving that it is composite

OpenStudy (fibonaccichick666):

https://primes.utm.edu/notes/proofs/Theorem2.html I was thinking this in reverse

OpenStudy (fibonaccichick666):

well, I thought it was implied by the corollary

ganeshie8 (ganeshie8):

Ahh that works in one direction i think.. \(a^n-1\) is prime \(\implies\) \(n\) is prime and \(a=2\)

OpenStudy (fibonaccichick666):

but then why not the other way? I mean this number ends in 7

ganeshie8 (ganeshie8):

yes units place is 7

OpenStudy (fibonaccichick666):

it is congruent to 3 mod 4 I expect

OpenStudy (fibonaccichick666):

but unless we rewrite that exponent somehow...

OpenStudy (fibonaccichick666):

ok

OpenStudy (fibonaccichick666):

so if we were to say \[(2^2)^{5903/2}\]

ganeshie8 (ganeshie8):

interesting, yes it is 3 mod 4

OpenStudy (danjs):

is that one of those mersene numnbers

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

it is a mersenne number but not a mersenne prime

OpenStudy (danjs):

saw a Numberphile video on those a while back, pretty interesting

OpenStudy (fibonaccichick666):

so, it's not a prime, but it is congruent to 3 mod 4.... hmmmmm

ganeshie8 (ganeshie8):

it is composite for sure you an take my word for it till we find a divisor :)

OpenStudy (fibonaccichick666):

ok just something I noticed, doesn't the number of prime factors for these numbers show the layout of twin primes???? http://mathworld.wolfram.com/MersenneNumber.html

OpenStudy (fibonaccichick666):

it's super cool

OpenStudy (fibonaccichick666):

....if not, I know I've seen that sequence before....

OpenStudy (fibonaccichick666):

maybe it's the number of numbers between primes, but I know I have seen that sequence before... I used to write out a bunch of stuff like that

OpenStudy (fibonaccichick666):

yay! a new project again!

ganeshie8 (ganeshie8):

i checked those oeis lists but couldnt find any of the divisors here is one prime divisor of 2^5903 - 1 : `15489473` i have one more prime divisor thats much smaller than this i really don't know how many more prime divisors are there as the number is pretty large its getting pretty late over here.. il leave this question open and checkout in the morning. good luck everyone :)

OpenStudy (fibonaccichick666):

There should be quite a few http://oeis.org/A046051 It is the 777th Mersenne number, but IDK if the prime factorizations have been done to that high

ganeshie8 (ganeshie8):

how do i expand the list ?

OpenStudy (fibonaccichick666):

that I am not sure of yet

ganeshie8 (ganeshie8):

i mean that link is showing only the first few numbers

ganeshie8 (ganeshie8):

how did you figure out it is 777th mersenne number ?

OpenStudy (fibonaccichick666):

5903 is the 777th prime, the numbers are generated by n where n maps p_1-->2 p_2-->3 so on and so on, right?

ganeshie8 (ganeshie8):

Omg yess got you :)

OpenStudy (fibonaccichick666):

so if someone else has charted that, we could know how many factors there are easily, I would hope

OpenStudy (fibonaccichick666):

so 777 is composite does imply that your number is composite quite quickly, I wonder if we said the number of factors of yours are \(Number~ of ~factors(M_3*M_7*M_37)\) if that would hold... That would give you 2 prime factors.

OpenStudy (fibonaccichick666):

so if you have one, there is only one more?

OpenStudy (fibonaccichick666):

nope... That unfortunately was disproven I think

OpenStudy (fibonaccichick666):

but hey, I broke a website for ya, so be happy :)

OpenStudy (fibonaccichick666):

9551342188655138465520630850748050237740825070468892272952870924644971060885980014797101592975604225394232178584042095438080874653033000331395103467735281328601695054847936491311527786671682592162023203313739961799537721188890561032015232740739319738333112751236122135697776461411064734226086042781330708005573085444012640234119134025335802746712411549041279893254970691479356955432922200407909731340402070631008165402702077757381508332714274047231528316463577958424135101308723835511153839289407331387837071445007284264743628515026257087054681121593156860426422949386744082801723391376251960500104779889096896085796997694877959326790879529168673362194818833716690707451436228868264143311645972611755728082963833538685570211392747634090165642002248970512159936093872616529809897187324891431584983442345523130705081126977817771847545510095508799677682071192057413296958123243781136260300990024753753003485045558304750628504680559293456024198635220012721581300440268692644730888654278822309815992095786163321036302624700675030450740050850140814899538810965000023356460637220447650815182582204557020578760952164279444934780918311159828233237536134274307033525441656137713203494335485237821182337824749946436459089651152619038477420024751551968999810256344722627877341947449539212500064909603404791421998814974051217774484619860053572148942068366595513913435156942402630948983804055905651162213754331132351212284929810561393370511275820187878030146314727136735033789567398009005283583734875313005061987943067150159176688822310958217907933010033542184922276062838408401300536480082515947239701583322577589591154094907178711808300932326795024101385254249838170624555401802817597058566015839803846142809468260749641012050369357086530748854455111175076440017518292017809047571302187007 That is our number

OpenStudy (princeharryyy):

3 is the answer @ganeshie8 (a^p -1)/p is fully divisible by p where a and p are prime numbers... fermat's remainder theorem.

OpenStudy (fibonaccichick666):

a^p-1/p has a remainder of 1 by fermat's little thm, no?

OpenStudy (anonymous):

ermm phi function failed :P i'll try in morning

OpenStudy (anonymous):

:3 i wish 5903 is not prime xD

OpenStudy (jhannybean):

Ohhh, @Kainui posted a really cool prime factorization link in one of his threads. I'll see if I can find it.

OpenStudy (jhannybean):

Might help? Maybe.

OpenStudy (anonymous):

first solution 11087

OpenStudy (anonymous):

@Marki \[ 3231\equiv 2^{5903}-1\pmod{11087} \]

OpenStudy (anonymous):

typo i meant 11807

OpenStudy (anonymous):

Hmmm, that looks like what I'm getting: http://jsfiddle.net/wio_dude/c0cj428m/

OpenStudy (anonymous):

I had tried earlier but wasn't finding any results because I had my exponent incorrect.

ganeshie8 (ganeshie8):

Wow! looks you guys nailed it ! wio that program runs super fast xD

ganeshie8 (ganeshie8):

@Marki im sure you must be having some cool number theory trick for finding divosors - please post your method when free (:

OpenStudy (kainui):

My method broke but let me type it out since I think it's interesting.

ganeshie8 (ganeshie8):

yes kai xD

OpenStudy (kainui):

\[\Large \sum_{k=0}^{5904} \frac{5904!}{k! \ (5904 - k)!}=2^{5904}\] Start here and take out the first and last terms of the series \[\Large \frac{1}{2} \sum_{k=1}^{5903} \frac{5904!}{k! \ (5904 - k)!}=2^{5903}-1\] But after typing this I just realized, the summation is symmetric, so we can just write it as: \[\Large \frac{5904! }{2(2952!)^2} + \sum_{k=1}^{2951} \frac{5904!}{k! \ (5904 - k)!}=2^{5903}-1\] So it looks like we can probably factor out something here, but I don't know what quite yet.

ganeshie8 (ganeshie8):

nice :) basically your strategy translates the problem to one of finding a common factor among first 5904 binomial coefficients very interesting idea XD il need to think about it more

OpenStudy (kainui):

I think "strategy" is too nice, I'm really just randomly trying stuff. I did come across this nice formula though: \[\Large \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) = \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) \frac{a+1}{2(a-k)}=2^a-1\]

OpenStudy (kainui):

It almost seems wrong when you look at it, but it is actually right. Haha

ganeshie8 (ganeshie8):

left and right sides are familiar to me middle part looks new.. leme think a bit..

OpenStudy (anonymous):

Hey kai, how did you get \[\Large \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) = \color{red}{\sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) \frac{a+1}{2(a-k)}=2^a-1}\] everything in red

OpenStudy (kainui):

I think @ganeshie8 is trying to figure it out first. =P

OpenStudy (anonymous):

Oh I see, ok

OpenStudy (kainui):

Or maybe I did write it wrong #_# let me check it... XD Sorry haha

ganeshie8 (ganeshie8):

right most reddy thing is a result of number of different ways of choosing a subset from a given set middle part doesnt belong there

ganeshie8 (ganeshie8):

\[\sum \limits_{k=0}^n \binom{n}{k} = 2^n\]

ganeshie8 (ganeshie8):

this is an identity, im still trying to figure out middle part..

OpenStudy (kainui):

Corrected, whoops\[\Large \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) = \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) \frac{a+1}{2(a+1-k)}=2^a-1\]

OpenStudy (kainui):

Start with binomial theorem\[\Large \sum_{k=0}^{a+1} \frac{(a+1)!}{k! \ (a+1-k)!}=2^{a+1}\] Pull out the first and last terms (they're both 1) and inside pull out some numbers from the factorials \[\Large 2+\sum_{k=1}^{a} \frac{a!}{k! \ (a-k)!} \frac{a+1}{a+1-k}=2^{a+1}\] Move the 2 to the other side and factor out a 2 \[\Large \frac{1}{2} \sum_{k=1}^{a} \frac{a!}{k! \ (a-k)!} \frac{a+1}{a+1-k}=2^a-1\] There we go. =P

OpenStudy (kainui):

I wonder if we can rearrange that second term in there to be like the geometric series and sort of do some kind of cauchy product with it.

ganeshie8 (ganeshie8):

Looks nice :) \[\binom{n}{k} = \binom{n}{k-1}\frac{n-k+1}{k}\]

ganeshie8 (ganeshie8):

that looks like a special case of newton's identity \[\binom{n}{k} \binom{k}{r} = \binom{n}{r} \binom{n-r}{k-r}\] lets see if we can use all these and pull some factor out xD

OpenStudy (kainui):

If we can do it generally, we would have a better idea of which mersenne numbers are primes maybe. Or maybe that will actually prevent us from doing it generally in the first place haha.

ganeshie8 (ganeshie8):

\[\frac{a+1}{a+1-k} = \frac{1}{1 - \frac{k}{a+1}}\] like this kai ?

OpenStudy (kainui):

Yeah that's what I meant earlier. However that's inside a sum rather than outside it, so it's sort of not going to work I think.

OpenStudy (kainui):

Playing around I got this: \[\Large \frac{1}{2} \sum_{n=0}^\infty \sum_{k=1}^a \left(\begin{matrix}a \\ k\end{matrix}\right) \left( \frac{k}{a+1} \right)^n=2^a-1\]

OpenStudy (kainui):

More general alternative forms I'm playing with are: a odd: \[\Large \sum_{k=1}^{(a-1)/2} \left(\begin{matrix}a \\ k\end{matrix}\right) \frac{a+1}{a+1-k}=2^a-1\] a even: \[\Large \frac{a!}{2 (a/2)!^2} + \sum_{k=1}^{a/2} \left(\begin{matrix}a \\ k\end{matrix}\right) \frac{a+1}{a+1-k}=2^a-1\]

OpenStudy (kainui):

I am getting tired so maybe I messed one of those up, but I know this is true: \[\Large \frac{5904! }{2(2952!)^2} + \sum_{k=1}^{2951} \frac{5904!}{k! \ (5904 - k)!}=2^{5903}-1\] And upon looking at it, we can easily factor out 5904 and leave 5903! behind since that will always leave integers, so we're safe there. Now we just need to see what factors of 5904 we can factor out of the other term. Well obviously since 2^asdofuao-1 is odd we just divide 5904 by 2 until we get an odd number which is 369. We factor this tiny number into 3*3*41. Since I know from mods that the last digit of the number is 7, 41 is not a factor, so I pray that it's 3 and post this and go to sleep.

OpenStudy (anonymous):

>.< i was gonna post something like this

OpenStudy (anonymous):